Status Report

ISS On-Orbit Status 5 Jul 2002

By SpaceRef Editor
July 5, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previou
sly.  A day of medical tests for the crew.

Early before breakfast, the Russian medical experiment PZE MO-11 (blood che
mistry analysis) completed by CDR Korzun and FE-2 Treschev, each taking turn
s in assisting as CMO (crew medical officer).  For the test, the crew y
esterday had destowed kits and accessories of the German-built Reflotron 4 b
lood analyzer (earlier versions were already on Mir).  The analysis pro
cedure was documented with the Nikon F5 still camera.  [Today, fresh
blood was drawn from the subject’s finger with a pipette, after the subject
had imbibed 250 ml of warm water or plain (unsugared) tea.  Clinical p
arameters were then determined from the collected samples.  Using vario
us reagent tabs, the blood was tested for such parameters as hemoglobin, glu
cose, bilirubin, amylase, uric acid, triglycerides, urea, etc.  The tub
es with blood samples were then stored in  the refrigerator.  The
Reflotron 4 uses 40 W of power, supplied by the SM electrical system.]

Later, the U.S. PHS (periodic health status) evaluation, without blood
labs and MO-9 urinalysis, was performed on all crewmembers. First FE-1 Peggy
Whitson acted as CMO for Korzun and Treschev, then Korzun assisted Whitson
with her own test.  [The PHS, guided by software on the MEC (medical
equipment computer), is performed once every 30 days for each crewmember an
d two weeks before landing and as clinically indicated.  The PHS evalua
tions alternate between the one with blood labs and the one without.  T
he tests are documented afterwards, and physical exercise should then follow

Peggy Whitson spent some time familiarizing herself with the SUBSA expe
riment (Solidification Using a Baffle in Sealed Ampoules), then installed th
e SUBSA thermal chamber and power control module (PCM) in the work volume (W
V) of the new Microgravity Science Glovebox (MSG). Since MSG is not commissi
oned yet for operations, the rack was not activated and was without power or
lighting.  Peggy only installed the hardware and its cabling inside th
e WV, documented by Treschev photographically for historical record.  T
he setup will be completed at a later date.

Peggy then took video of the SAMS (space acceleration measuring system) sen
sor enclosure (SE), capturing the orientation of the SAMS SE deployed during
the SUBSA setup.

FE-2 Sergei Treschev performed repair work on the Service Module (SM) onboa
rd measuring system (SBI), replacing a T282 thermal sensor with a new unit.
 Two hours were allocated for the task, which included soldering connec
ting wires and continuity testing of the sensor contacts and the BKS onboard
cable network.  The work was documented with the Kodak 760 digital sti
ll camera (DSC) for inspection by ground specialists.

CDR Valery Korzun used the DSC to take pictures of the internal surface of
the docking assembly used for the Progress linkup (which had been moved out
of the hatchway after the docking).  The objective was to record images
the scratch mark made by the head of the docking probe on the internal surf
ace of the cone, to be used for improving existing characterization of docki
ng parameters.  Later in the day (11:30 am EDT), Korzun downlinked the

The two cosmonauts continued unloading, transferring and stowing Progress c
argo, with IMS (inventory management system) tracking.

FE-1 Whitson performed the T + 2 days microbiological analysis of water sam
ples with the WMK (water microbiology kit).

Treschev transferred stored exercise data from the TVIS (treadmill), RED (r
esistive device) and CEVIS (cycle) to the MEC (medical equipment computer),
for later downlinking to the ground.

Daily routine tasks were completed by CDR Korzun (inspection of BRPK-2 wate
r condensate separator, and IMS delta file preparation), FE-2 Treschev (SOSh
life support system servicing), and FE-1 Whitson (Lab payloads status check

The crewmembers also performed the daily physical exercise routine.

Troubleshooting by the ground continues on the CSA-CP (compound specific an
alyzer-combustion products) sampling pump, which is not functioning properly
when Whitson wants to use it with certain filters.

Yesterday, Peggy videotaped a "guided tour" of the ISS without au
dio on PAO request per "job jar", to be played back in a live PAO
event with the CNN network on 7/9, with the crew then narrating over the vid
eo live in real-time.

Today’s targets of the U.S. crew earth observation (CEO) program were Ma
(this target is difficult to acquire at this time of the year,
but a break in the weather was possible this pass. The crew was asked to us
e the long lenses of the ESC [electronic still camera] to map the details of
the coral reef formations along the east coast of the Malay Peninsula)
Cloud Vortices (Dynamic Event Site: Brisk northerly winds were dr
iving a large field of stratus clouds past the Madeira and Canary Island arc
hipelagos this pass setting up favorable conditions for cloud vortices downw
ind. Of interest: looking to the right of track for these features where the
mountaintops of the islands rise above the surrounding cloud field and disr
upt the airflow)
, Angolan Biomass Burning (numerous large fire
s are now being detected by satellite over much of southern Africa. As the I
SS tracked southeastward over southern Zaire and northeastern Angola, the cr
ew was to look to the right of track to detect fire plumes and their point s
, Congo-Zimbabwe Biomass Burning (after Zaire and Angol
a the station tracked SE over Zambia and Zimbabwe.  Of interest for thi
s target: fire plumes and their sources)
, Canadian Rocky Mountains (the weather was questionable for the SW half of this target area. &nbs
p;Crew was to try for oblique, context views to the left of track and may al
so have detected forest fires reported further north)
, and Eastern Un
ited States
(if the ongoing smog conditions persisted, the crew was a
dvised to try for oblique and limb views to the right of track this pass to
document the thickness and extent of aerosols over the upper Ohio Valley and
Mid-Atlantic States)

CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 1:25 pm EDT):

Environmental Control and Life Support (ECLSS) and Thermal
Control (TCS):

Elektron O2 generator is powered On (32-amp mode). Vozdukh CO2 s
crubber is ON in MANUAL cycle mode #5 (vacuum pump failed).  U.S. CDRA
CO2 scrubber is Off (failed).  BMP Harmful Impurities unit: Absorbent b
ed #1 in Purify mode, bed #2 in Purify mode.

SM Working Compartment:  Pressure (mmHg) — 757, temperature (d
eg C) — 26.0, ppO2 (mmHg) — 156.5, ppCO2 (mmHg) — 2.9.

SM Transfer Compartment:  Pressure (mmHg) — 762, temperature (
deg C) — 21.0.

FGB Cabin:  Pressure (mmHg) — 756, temperature (deg C) — 20.0

Node:  Pressure (mmHg) — 748.98, temperature (deg C) — 22.2 (
shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.

U.S. Lab:  Pressure (mmHg) — 750.88 temperature (deg C) — 24.
8, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;

Joint Airlock (Equip. Lock):  Pressure (mmHg) — 750.67, temper
ature (deg C) — 22.7; shell heater temp (deg C) — 23.5, ppO2 (mmHg) — n/a
; ppCO2 (mmHg) — n/a.

PMA-1:  Shell heater temp (deg C) — 22.7

PMA-2:  Shell heater temp (deg C) — 18.9

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not avail
able because MCA [major constituent analyzer] is failed and in Extended Life
[3D a state that preserves mass spectrometer vacuum but produces
no pp data]). MSA (mass spectrometer assembly) and VGA (verification gas as
sembly) were replaced, but some more work needs to be done.

Electrical Power Systems (EPS):

Beta Gimbal Assembly (BGA) 2B in Autotrack mode; BGA 4B in Directed
position (non solar tracking).

SM batteries: (no data available).

FGB: Battery #6 is offline (failed); all other batteries (5) are in 93Partia
l charge94 mode.

Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:

Air conditioner SKV-1 is Off. SKV-2 is On.

Command & Data Handling Systems:

C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1 is in s

GNC-1 MDM is prime; GNC-2 is Backup.

LA-1, LA-2 and LA-3 MDMs are all operating.

PL-1 MDM is operational; PL-2 MDM is Off.

APS-1 (automated payload switch #1) and APS-2 are both On.

SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational.

SM Central Computer (TsVM): 1 lane (of 3) operational.

Attitude Source:

3 CMGs online.

State vector — US GPS (SIGI string 1)

Attitude — Russian segment

Angular rates — US RGA1 (rate gyro assembly 1)

Communications Systems:

All Russian communications & tracking systems are nominal.

S-band is operating nominally.

Ku-band is operating nominally.

Audio subsystem operating nominally.

Video subsystem operating nominally.

MCOR (medium-rate communications outage recorder) was R&R’d today and i
s operating nominally.


SSRMS/Canadarm2 at MBS PDGF (mobile base system/power & data gr
apple fixture) and Lab PDGF, with Keep Alive power on both strings (based on

MBS:  Operational on Redundant string, keep-alive power on Prime strin

RWS (robotics workstations): Cupola is Active, Lab RWS is Off.

ISS Orbit (as of this morning, 7:35 am EDT [3D epoch]):

Mean altitude — 391.6 km

Apogee — 396.8 km

Perigee — 386.4 km

Period — 92.4 min.

Inclination (to Equator) — 51.64 deg

Eccentricity — 0.0007711

Orbits per 24-hr. day — 15.58

Altitude decrease — 110 m (mean) in last 24 hours

Revolutions since FGB/Zarya launch (Nov. ’98) — 20688

Current Flight Attitude — LVLH (local vertical/local horizontal 3D”earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg,
pitch: 7.8 deg, roll: 0 deg])

For more on ISS orbit and naked-eye visibility dates/times, see

SpaceRef staff editor.